# DSCA: A Digital Subtraction Angiography Sequence Dataset and Spatio-Temporal Model for Cerebral Artery Segmentation

### v1

@@ IV-C Cerebrovascular Segmentation Results
TEB TF STF BV MTA BV MTA BV MTA BV MTA BV MTA
76.76±$\pm{0.45}$ 82.54±$\pm{2.42}$ 86.45±$\pm{0.31}$ 86.22±$\pm{2.50}$ 85.29±$\pm{0.45}$ 85.70±$\pm{2.50}$ 88.06±$\pm{0.59}$ 88.08±$\pm{2.20}$ 92.41±$\pm{0.22}$ 89.87±$\pm{2.51}$
77.40±$\pm{0.26}$ 84.02±$\pm{4.82}$ 86.87±$\pm{0.18}$ 88.25±$\pm{4.87}$ 86.12±$\pm{0.44}$ 87.20±$\pm{5.12}$ 87.99±$\pm{0.67}$ 91.03±$\pm{3.75}$ 92.83±$\pm{0.21}$ 91.30±$\pm{4.67}$
78.02±$\pm{0.22}$ 86.00±$\pm{1.43}$ 87.25±$\pm{1.43}$ 89.91±$\pm{1.52}$ 86.56±$\pm{0.10}$ 91.86±$\mathbf{\pm{1.05}}$ 88.30±$\pm{0.22}$ 92.81±$\pm{1.32}$ 93.05±$\pm{0.05}$ 93.18±$\pm{1.19}$
77.83±$\pm{0.33}$ 86.14±$\pm{2.21}$ 87.14±$\pm{0.21}$ 89.99±$\pm{2.29}$ 86.84±$\mathbf{\pm{0.60}}$ 88.99±$\pm{2.18}$ 87.78±$\pm{0.66}$ 93.03±$\pm{2.63}$ 93.18±$\mathbf{\pm{0.29}}$ 93.34±$\pm{2.00}$
78.09±$\mathbf{\pm{0.20}}$ 88.22±$\mathbf{\pm{1.52}}$ 87.32±$\mathbf{\pm{0.13}}$ 92.26±$\mathbf{\pm{1.56}}$ 86.40±$\pm{0.39}$ 91.37±$\pm{1.57}$ 88.59±$\mathbf{\pm{0.53}}$ 94.01±$\mathbf{\pm{1.40}}$ 92.98±$\pm{0.18}$ 94.98±$\mathbf{\pm{1.39}}$
@@ IV-D Ablation Study
We performed extensive ablation experiments on the DSCA dataset to illustrate the effectiveness of the components in DSANet. Utilizing nnUNet as the baseline, we systematically evaluated the influence of each component on the segmentation results, as shown in Table IV.
IV-D1 Ablation for temporal encoding branch
In this section, we present a comparative analysis of metrics including JAC, Dice, SEN, PRE, and AUC to assess the impact of integrating TEB. We incorporate TEB with nnUNet while maintaining consistent experimental settings. The ablation results, presented in the 4th row of Table IV, demonstrate a significant improvement in both BV and MTA segmentation. This is primarily attributed to the integration of dynamic flow information. Notably, the incorporation of TEB resulted in an increase of over 1% in JAC, Dice, SEN, PRE, and AUC in MTA.
IV-D2 Ablation for TemporalFormer
Subsequently, we validate the impact of TF based on the established foundation of the TEB. As depicted in the fifth row of Table IV, all experimental outcomes have shown improvement to varying degrees compared to the fourth row. The enhancement in the segmentation results for MTA is more prominent than that for BV. Specifically, the SEN indicator has reached its optimal level in MTA in this method, reaching 91.86%.
IV-D3 Ablation for spatio-temporal fusion module
Similarly, we assess the superiority of the STF module in conjunction with the TEB. Comparing the results in the sixth and fourth rows of Table IV, it is obvious that incorporating the STF module promotes better segmentation performance. Among them, the SEN and AUC metrics reach their optimal BV segmentation for our method, increasing the TEB by 0.72% and 0.35%, respectively. In the PRE metric, it slightly drops by 0.28% and 0.21% in BV, while outperforming them by 4.95% and 2.00% in MTA, compared to the baseline and the TEB, respectively. The last row in Table IV outlines our results. A comprehensive review of the ablation experiments in the table explicitly illustrates the effectiveness of our method and its components.
@@ IV-D3 Ablation for spatio-temporal fusion module
Figure 7: Several instances where the MTA class is employed to assess stenosis. The more intense the red color, the more severe the stenosis.
@@ V Discussion
V-A DSCA: Promoting Clinical Cerebrovascular Analysis
The DSCA dataset comprises 224 sequences, each meticulously annotated pixel by pixel using MIP images. It includes ICA, ECA, and VA, with further subclassification of ICA and VA into BV and MTA. This classification closely aligns with the requirements of clinical analysis. Previous studies have demonstrated the value of precisely measuring carotid stenosis [37]. The extraction of MTA facilitates automated diameter calculations for ICA and VA, particularly in evaluating biomarkers at stenosis sites. Fig. 7 reveals several instances where MTA is used to assess stenosis in the main trunk of the artery. In Fig. 7-(a), (b), (d), the diameter of the stenosis region can be easily determined from the segmented MTA. Moreover, the color representation of the arterial main trunk appears normal in the absence of stenosis, as illustrated in Fig. 7-(c), (e). Similarly, vascular density can be obtained for individually extracted BV to support clinical applications effectively, as demonstrated in the previous research for evaluating thrombolysis scores in ischemic stroke [29].
V-B Why Resample 8 Frames for Training?
@@
V-C DSANet: Effectively Handle Complex Situations
It is challenging to accurately segment CA due to small vessels with low visibility and ambiguity between vessels and residual skull structures. To further validate the effectiveness of our approach, we generate heatmaps for the final layer output of the decoder and make comparisons with state-of-the-art methods, as shown in Fig. 10. In the 2nd and 4th rows, all methods exhibit a positive response in MTA. Additionally, a comparison of the 1st and 3rd rows reveals that our DSANet produces a stronger response than UNet, TransUNet, and nnUNet on small vascular structures in BV, particularly in the areas highlighted by the white circles. Meanwhile, UNet and TransUNet are more susceptible to background artifacts and noise than the proposed DSANet, as shown in the 1st and 2nd columns of BV. This highlights the superior performance of our method in suppressing background interferences.
@@ VI Conclusion and Future Works
In this paper, we have established the DSCA dataset, a large, carefully designed, and systematically annotated dataset for CA segmentation in DSA sequences. To our knowledge, the DSCA is the first publicly available DSA sequence dataset. It not only has pixel-wise vessel annotations but also includes subdivisions for BV and MTA, which are closely related to the clinical treatment of CVDs. This dataset will be highly beneficial for researchers in CVD studies. Furthermore, we have proposed DSANet, a sptaio-temporal network for CA segmentation. Unlike 2D segmentation methods, DSANet incorporates an independent temporal encoding branch to capture dynamic temporal information. It integrates TemporalFormer and Spatio-Temporal Fusion modules to further process spatio-temporal features. Extensive experiments demonstrate the effectiveness of our method, particularly small vessels, holding significant potential for clinical applications.

### v2

@@ IV-C Cerebrovascular Segmentation Results
TEB TF STF BV MAT BV MAT BV MAT BV MAT BV MAT
MinIP 76.76 $\pm$ 0.45 82.54 $\pm$ 2.42 86.45 $\pm$ 0.31 86.22 $\pm$ 2.50 85.29 $\pm$ 0.45 85.70 $\pm$ 2.50 88.06 $\pm$ 0.59 88.08 $\pm$ 2.20 92.41 $\pm$ 0.22 89.87 $\pm$ 2.51
Sequence 76.77 $\pm$ 0.40 84.91 $\pm$ 2.23 86.45 $\pm$ 0.27 89.11 $\pm$ 2.31 85.12 $\pm$ 0.30 88.00 $\pm$ 2.04 88.29 $\pm$ 0.58 91.87 $\pm$ 2.51 92.33 $\pm$ 0.15 92.15 $\pm$ 2.31
Sequence + MinIP 77.52 $\pm$ 0.22 85.90 $\pm$ 2.10 86.94 $\pm$ 0.17 89.88 $\pm$ 2.29 85.73 $\pm$ 0.32 88.76 $\pm$ 2.09 88.56 $\pm$ 0.31 91.93 $\pm$ 2.57 92.64 $\pm$ 0.15 93.00 $\pm$ 1.88
77.40 $\pm$ 0.26 84.02 $\pm$ 4.82 86.87 $\pm$ 0.18 88.25 $\pm$ 4.87 86.12 $\pm$ 0.44 87.20 $\pm$ 5.12 87.99 $\pm$ 0.67 91.03 $\pm$ 3.75 92.83 $\pm$ 0.21 91.30 $\pm$ 4.67
78.02 $\pm$ 0.22 86.00 $\pm$ 1.43 87.25 $\pm$ 1.43 89.91 $\pm$ 1.52 86.56 $\pm$ 0.10 91.86 $\pm$ 1.05 88.30 $\pm$ 0.22 92.81 $\pm$ 1.32 93.05 $\pm$ 0.05 93.18 $\pm$ 1.19
77.83 $\pm$ 0.33 86.14 $\pm$ 2.21 87.14 $\pm$ 0.21 89.99 $\pm$ 2.29 86.84 $\pm$ 0.60 88.99 $\pm$ 2.18 87.78 $\pm$ 0.66 93.03 $\pm$ 2.63 93.18 $\pm$ 0.29 93.34 $\pm$ 2.00
78.09 $\pm$ 0.20 88.22 $\pm$ 1.52 87.32 $\pm$ 0.13 92.26 $\pm$ 1.56 86.40 $\pm$ 0.39 91.37 $\pm$ 1.57 88.59 $\pm$ 0.53 94.01 $\pm$ 1.40 92.98 $\pm$ 0.18 94.98 $\pm$ 1.39
TABLE VI: Ablation study of these components of the proposed DSANet for the entire vessel. The results are average±$\pm$±standard deviation obtained from the five-fold cross-validation experiment. All metrics are expressed as percentages (%).
Ablation Sets clDice JAC Dice SEN PRE AUC
TEB TF STF
MinIP 87.54 $\pm$ 0.25 81.70 $\pm$ 0.23 89.69 $\pm$ 0.14 88.45 $\pm$ 0.28 91.17 $\pm$ 0.17 94.01 $\pm$ 0.14
Sequence 87.78 $\pm$ 0.19 81.67 $\pm$ 0.23 89.65 $\pm$ 0.14 88.21 $\pm$ 0.16 91.40 $\pm$ 0.27 93.89 $\pm$ 0.08
Sequence + MinIP 88.25 $\pm$ 0.15 82.23 $\pm$ 0.16 90.07 $\pm$ 0.10 88.65 $\pm$ 0.32 91.72 $\pm$ 0.31 94.06 $\pm$ 0.15
88.29 $\pm$ 0.11 82.39 $\pm$ 0.11 90.11 $\pm$ 0.07 89.13 $\pm$ 0.37 91.30 $\pm$ 0.32 94.34 $\pm$ 0.17
88.42 $\pm$ 0.13 82.60 $\pm$ 0.15 90.19 $\pm$ 0.09 89.28 $\pm$ 0.10 91.43 $\pm$ 0.25 94.40 $\pm$ 0.05
88.41 $\pm$ 0.18 82.56 $\pm$ 0.24 90.17 $\pm$ 0.14 89.60 $\pm$ 0.34 91.12 $\pm$ 0.39 94.38 $\pm$ 0.16
88.58 $\pm$ 0.13 82.76 $\pm$ 0.16 90.33 $\pm$ 0.09 89.30 $\pm$ 0.23 91.56 $\pm$ 0.32 94.44 $\pm$ 0.11
@@ IV-D Ablation Study
We performed extensive ablation experiments on the DSCA dataset to illustrate the effectiveness of the components in DSANet. Utilizing nnUNet as the baseline, we systematically evaluated the influence of each component.
IV-D1 Ablation for MinIP and sequence
In Table V and VI, the ablation experiments for MinIP and sequence are performed. As shown from the third to the fifth row of TABLE V and VI, there are slight differences in performance between MinIP input and sequence input. While the sequence input does not bring additional spatial information, it introduces temporal information. MinIP, although not adding extra information, contains features that are easier to extract than the sequence input. Using both MinIP and sequence inputs together results in some performance improvement. Interestingly, for MAT, sequence input performs better than MinIP input, suggesting that the introduction of dynamic temporal information helps in distinguishing between MAT and BV more effectively.
IV-D2 Ablation for temporal encoding branch
In this section, we present a comparative analysis of metrics including JAC, Dice, SEN, PRE, and AUC to assess the impact of integrating TEB. We incorporate TEB with nnUNet while maintaining consistent experimental settings. The ablation results, presented in the 4th row of Table V, demonstrate a significant improvement in both BV and MAT segmentation. This is primarily attributed to the integration of dynamic flow information. Notably, the incorporation of TEB resulted in an increase of over 1% in JAC, Dice, SEN, PRE, and AUC in MAT.
IV-D3 Ablation for TemporalFormer
Subsequently, we validate the impact of TF based on the established foundation of the TEB. As depicted in the fifth row of Table V, all experimental outcomes have shown improvement to varying degrees compared to the fourth row. The enhancement in the segmentation results for MAT is more prominent than that for BV. Specifically, the SEN indicator has reached its optimal level in MAT in this method, reaching 91.86%.
IV-D4 Ablation for spatio-temporal fusion module
Similarly, we assess the superiority of the STF module in conjunction with the TEB. Comparing the results in the sixth and fourth rows of Table V, it is obvious that incorporating the STF module promotes better segmentation performance. Among them, the SEN and AUC metrics reach their optimal BV segmentation for our method, increasing the TEB by 0.72% and 0.35%, respectively. In the PRE metric, it slightly drops by 0.28% and 0.21% in BV, while outperforming them by 4.95% and 2.00% in MAT, compared to the baseline and the TEB, respectively. The last row in Table V outlines our results. A comprehensive review of the ablation experiments in the table explicitly illustrates the effectiveness of our method and its components. After adding both the TF and SFT components, our method falls behind the methods that only add one component in terms of the SEN metric. However, DSANet demonstrates superior performance in other metrics, such as the PRE metric. This suggests that adding one component increases the detection of false positives in the background. By combining these two components, our method is able to achieve better overall performance. As shown in Table VI, these are the ablation study results for the entire vascular structure. After incorporating the sequence, both our method and nnUNet exhibited some performance improvement. However, after adding the TF and STF components, the performance difference was not very significant. We think this may be because the model takes into account the distinction between BV and MAT during training, leading to only a slight improvement in the overall vascular segmentation metrics.
@@ IV-D4 Ablation for spatio-temporal fusion module
TABLE VII: Experiments for TF in parallel, cldice loss, and variable length sequence. ALL represents the entire vessel.
Experiment Sets JAC Dice AUC
BV MAT ALL BV MAT ALL BV MAT ALL
Parallel 77.89 $\pm$ 0.28 87.21 $\pm$ 1.90 82.78 $\pm$ 0.11 87.14 $\pm$ 0.22 90.93 $\pm$ 1.98 90.34 $\pm$ 0.07 92.99 $\pm$ 0.10 94.54 $\pm$ 1.49 94.45 $\pm$ 0.10
clDice 66.04 $\pm$ 0.81 75.45 $\pm$ 1.34 74.65 $\pm$ 0.48 78.75 $\pm$ 0.58 81.32 $\pm$ 1.71 85.11 $\pm$ 0.32 87.80 $\pm$ 0.53 86.33 $\pm$ 1.20 90.68 $\pm$ 0.40
DSANet_vl 75.95 $\pm$ 0.30 78.28 $\pm$ 3.55 81.08 $\pm$ 0.32 85.78 $\pm$ 0.18 82.08 $\pm$ 3.73 89.25 $\pm$ 0.20 92.53 $\pm$ 0.30 86.33 $\pm$ 3.53 94.07 $\pm$ 0.16
DSANet_fl 75.05 $\pm$ 0.34 76.53 $\pm$ 3.68 80.53 $\pm$ 0.28 85.14 $\pm$ 0.25 80.44 $\pm$ 3.74 $8$$\pm$ 0.19 92.02 $\pm$ 0.20 84.68 $\pm$ 3.68 93.70 $\pm$ 0.19
DSANet 78.09 $\pm$ 0.20 88.22 $\pm$ 1.52 82.76 $\pm$ 0.16 87.32 $\pm$ 0.13 92.26 $\pm$ 1.56 90.33 $\pm$ 0.23 92.98 $\pm$ 0.18 94.98 $\pm$ 1.39 94.44 $\pm$ 0.11
CAVE_vl 69.12 $\pm$ 0.84 69.32 $\pm$ 2.25 76.50 $\pm$ 0.68 80.59 $\pm$ 0.89 74.46 $\pm$ 2.21 86.12 $\pm$ 0.60 88.68 $\pm$ 0.63 77.98 $\pm$ 2.03 91.44 $\pm$ 0.53
CAVE_fl 68.91 $\pm$ 3.00 73.11 $\pm$ 3.20 75.70 $\pm$ 2.30 80.87 $\pm$ 2.57 77.80 $\pm$ 3.21 85.78 $\pm$ 1.73 91.84 $\pm$ 2.74 80.83 $\pm$ 3.30 93.86 $\pm$ 2.17
CAVE 69.11 $\pm$ 2.80 79.09 $\pm$ 2.16 75.89 $\pm$ 2.07 81.12 $\pm$ 1.88 83.70 $\pm$ 2.15 85.98 $\pm$ 1.44 91.22 $\pm$ 2.21 86.84 $\pm$ 1.91 93.15 $\pm$ 1.75
  • ’vl’ and ’fl’ represent the variable length and fixed length, respectively.
TABLE VIII: Experiments on DIAS dataset.
Methods clDice Dice SEN
nnUNet [22] 70.20 $\pm$ 4.65 78.09 $\pm$ 3.17 76.69 $\pm$ 7.81
VSSNet [35] 71.23 $\pm$ 5.65 78.34 $\pm$ 3.92 77.27 $\pm$ 7.73
DSANet 71.38 $\pm$ 5.27 78.37 $\pm$ 3.61 77.88 $\pm$ 8.15
Figure 7: Segmentation results on the DIAS dataset. Green, red, and blue represent true positive, false positive, and false negative, respectively.
@@ V Discussion
V-A Additional Experimental Analysis
V-A1 Experiment for TF in parallel
We compared the parallel execution of temporal attention and spatial attention in the TF component. The third row of Table VII demonstrates that performing in series is relatively better than parallel execution, especially in distinguishing between MAT and BV. We think directly combining the enhanced feature from the temporal and spatial dimensions could result in negative effects if performed in parallel.
V-A2 Experiment for clDice loss
clDice loss is a variation of Dice loss that encourages the network to focus more on the central topology of vessels. However, this emphasis can have unintended effects on larger vessels, potentially reducing segmentation accuracy for these structures. We have added an experiment in the fourth row of Table VII where we replaced CE loss and Dice loss with CE loss and clDice loss, but the results were not promising.
V-A3 Experiment for variable-length sequence input
While using variable-length sequence input may seem advantageous, vascular segmentation tasks typically use fixed-length inputs or a fixed number of channels. This approach is also employed in DIAS [35], where variable-length sequences are resampled to fixed-length sequences. Moreover, variable-length sequence inputs restrict the batch size to 1, which is not conducive to model training. To verify the performance of raw sequence inputs, we have included an experiment with variable-length sequence input (DSANet_vl), as shown in Table VII. Since nnUNet does not support variable-length sequences, we developed a new training framework. For a fair comparison, we also implemented fixed-length training (DSANet_fl) by setting the batch size to 1. Similarly, we utilized CAVE to conduct the same training setting, resulting in both fixed-length (CAVE_fl) and variable-length (CAVE_vl) versions. Experimental results indicate that variable-length input outperforms fixed-length sequences, but is still behind the original DSANet.
V-A4 Experiment on the DIAS dataset
To further validate the generalization performance of our method, we conducted additional evaluations on the public DIAS dataset. To ensure a fair comparison, we retrained VSSNet on the DIAS dataset according to the setting in their paper. As presented in Table VIII, our DSANet demonstrates a slight performance improvement over both nnUNet and VSSNet. The qualitative results in Fig. 7 (highlighted in the orange box) illustrate that our method enhances cerebrovascular segmentation, underscoring its effectiveness and robustness in generalization.
Figure 8: Several instances where the MAT class is employed to assess stenosis. The more intense the red color, the more severe the stenosis.
V-B DSCA: Promoting Clinical Cerebrovascular Analysis
The DSCA dataset comprises 224 sequences, each meticulously annotated pixel by pixel using MinIP images. It includes ICA, ECA, and VA, with further subclassification of ICA and VA into BV and MAT. This classification closely aligns with the requirements of clinical analysis. Previous studies have demonstrated the value of precisely measuring carotid stenosis [46]. The extraction of MAT facilitates automated diameter calculations for ICA and VA, particularly in evaluating biomarkers at stenosis sites. Fig. 8 reveals several instances where MAT is used to assess stenosis in the main trunk of the artery. In Fig. 8-(a), (b), (d), the diameter of the stenosis region can be easily determined from the segmented MAT. Moreover, the color representation of the arterial main trunk appears normal in the absence of stenosis, as illustrated in Fig. 8-(c), (e). Furthermore, it is well known that the ICA and VA are divided into segments [47, 48]. Some researchers have studied the tortuosity [49, 50], morphometry [51], and diameter [52, 53] of the ICA and the clinical significance of specific ICA segments [54, 55]. Therefore, directly extracting the MAT is meaningful for these studies. Similarly, research on small vessels extends to studies on cerebral small vessel disease [56, 57] and cerebrovasculature labeling [33, 58], which remains a focus of ongoing clinical investigation. We believe our work can contribute to advancing these fields.
V-C Why Resample 8 Frames for Training?
@@
V-D DSANet: Effectively Handle Complex Situations
It is challenging to accurately segment CA due to small vessels with low visibility and ambiguity between vessels and residual skull structures. To further validate the effectiveness of our approach, we generate heatmaps for the final layer output of the decoder and make comparisons with state-of-the-art methods, as shown in Fig. 11. In the 2nd and 4th rows, all methods exhibit a positive response in MAT. Additionally, a comparison of the 1st and 3rd rows reveals that our DSANet produces a stronger response than UNet, TransUNet, and nnUNet on small vascular structures in BV, particularly in the areas highlighted by the white circles. Meanwhile, UNet and TransUNet are more susceptible to background artifacts and noise than the proposed DSANet, as shown in the 1st and 2nd columns of BV. This highlights the superior performance of our method in suppressing background interferences.
@@ VI Conclusion and Future Works
In this paper, we have established the DSCA dataset, a large, carefully designed, and systematically annotated dataset for CA segmentation in DSA sequences. To our knowledge, the DSCA is the first publicly available DSA sequence dataset. It not only has pixel-wise vessel annotations but also includes subdivisions for BV and MAT, which are closely related to the clinical treatment of CVDs. This dataset will be highly beneficial for researchers in CVD studies. Furthermore, we have proposed DSANet, a spatio-temporal network for CA segmentation. Unlike 2D segmentation methods, DSANet incorporates an independent temporal encoding branch to capture dynamic temporal information. It integrates TemporalFormer and Spatio-Temporal Fusion modules to further process spatio-temporal features. Extensive experiments demonstrate the effectiveness of our method, particularly small vessels, holding significant potential for clinical applications.